The demand for good R is QRd = 10, 000 − 4PR + 5PT + 2M + AR, where PR is the price of good R, PT is the price of good T, M is the buyers’ money income, and AR is the amount spent to advertise good R. Suppose PR = $50, PT = $100, M = $25,000, and AR = $1,000. The cross-price elasticity of goods R and T is:

  • −0.08.

  • None of the options.

  • −0.008.

  • 0.008.

  • 0.08.
Answer #1

Q = 10000-4*50+5*100+2*25000+1000 = 10000-200+500+50000+1000 = 61300

P = 100

coefficient of the good = 5

Cross price elasticity = 5*100/61300 = 0.008

option(C)